A hose can fill a swimming pool in 18 hours. Another hose needs 3 more hours to fill the pool than the two hoses combined. How long would it take the second hose to fill the pool?
Answer by edjones(8007)

1st hose does 1/18 the job in one hr.
The other hose needs x hours to do the job. So it can do 1/x of the job in one hour.
1/18 + 1/x = (18+x)/18x amount of the job done in one hr by the 2 hoses combined.
18x/(x+18) hours it takes for both hoses to fill the pool.
(18x/(x+18))+3 hours it takes the second hose to fill the pool (3 hrs more than the 2 hoses combined).
x=(18x/(x+18))+3
x-3=18x/(x+18)
(x-3)(x+18)=18x
x^2+15x-54=18x
x^2-3x-54=0
(x-9)(x+6)=0
x=9 hrs
.
Ed